OpenStudy (vianne):

What's 4.58333333333 in fraction form? I can't seem to find out. Too many 3's, and I don't have a specific calculator for it.

OpenStudy (anonymous):

http://www.webmath.com/dec2fract.html

OpenStudy (vianne):

4 - 27/-46? Could you explain that please?

OpenStudy (anonymous):

LOL

OpenStudy (vianne):

Simplifying our fraction = 4583333333/1000000000 = 4583333333/1000000000 = 4 583333333/1000000000 Seriously?

OpenStudy (anonymous):

:/

OpenStudy (zarkon):

is that the full number or do the 3's go on forever?

OpenStudy (vianne):

3's go on forever.

OpenStudy (vianne):

And I found the answer, 4 7/12

ganeshie8 (ganeshie8):

thats right! how did you get the answer ?

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

Some undiscovered magic?